• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 21:47:30 on February 22, 2011

    Entry number 348154

    keyword=Loop 1 and 2 Temperatures Stabilized for now

    The loop 1 and 2 temperatures have stabilize around 24.5 K and 23.4 K with
    a cryo flow around 7 g/s.  Not ideal, but not completely warm either. 
    I hope once Hall C finishes their cool down that our cyro flow will return
    to normal and the 15 K supply temperature will drop to normal as well 
    rather than the current 17 K.

    A copy of this log entry has been emailed to: kalyan,doug,jpchen